OpenStudy (anonymous):

Calculus: For 4x^(3)-3xy^(2)+y^(3)=28 find dy/dx at the point (3,4) on the curve. Can someone go step by step with me on how to do this please?

OpenStudy (anonymous):

yes easy.. first find the dy/dx of that equation.. bring all terms on one side.. you would get an expression in x and y, then subsititute 3,4 for x and y to get the final value!

OpenStudy (turingtest):

can you differentiate this implicitly, or are you having trouble with that?

OpenStudy (anonymous):

yeahh i think im having trouble with that; and okay but before i find the dy/dx, do i have to set it equal to zero or to y?

OpenStudy (turingtest):

when you take the derivative the right side will become zero, because the derivative of any constant is zero

OpenStudy (anonymous):

you cannot put x and y terms on separate sides, you need to directly differentiate.. !!

OpenStudy (anonymous):

quickly answer these what are the derivatives of 4x^3 3xy^2 y^3 ?? wrt x

OpenStudy (anonymous):

ohh okay! so then after finding the derivatives, do i plug in both the x and y values they gave me?

OpenStudy (anonymous):

12x^2

OpenStudy (anonymous):

3y^2

OpenStudy (anonymous):

and do i have to use the product rule for the second?

OpenStudy (turingtest):

no, that is the point, you need to use the chain rule here

OpenStudy (turingtest):

yes, and the product rule for the second

OpenStudy (turingtest):

so try again, what is the derivative of 3xy^2 with respect to x

OpenStudy (anonymous):

im sorry i actually haven't memorized the product rule yet :(

OpenStudy (turingtest):

well you sort of need to to do these. d/dx(f(x)g(x))=f(x)g'(x)+f'(x)g(x)

OpenStudy (anonymous):

http://www.math.ucdavis.edu/~kouba/CalcOneDIRECTORY/implicitdiffdirectory/ImplicitDiff.html Here are some good examples of implicite differentiation.

OpenStudy (anonymous):

okay before i do, since there is a 3, isn't that a constant so i don't have to include that?

OpenStudy (anonymous):

thank you eashmore!

OpenStudy (turingtest):

the 3 can stay outside if you want because for any constant C we have d/dx[Cf(x)]=Cf'(x) constants are funny that way

OpenStudy (anonymous):

is it 2xy+y^2?

OpenStudy (turingtest):

for what? xy^2 almost, but you didn't differentiate y implicitly. You need the chain rule: d/dx[f(g(x))]=f'(g(x))g'(x)

OpenStudy (anonymous):

what exactly does implicitly mean? and so do i use the chain rule for just the y^2 or the whole xy^3?

OpenStudy (binary3i):

whole

OpenStudy (anonymous):

derivative of y wrt x is dy/dx.. so if you have to find derivative of y^2 wrto x it is 2ydy/dx!

OpenStudy (turingtest):

when it is the derivative w/ respect to x, we need to differentiate y implicitly. Look what Mashy said. d/dx(y^2)=2yy', not just 2y

OpenStudy (turingtest):

that is the use of the chain rule. You need to use it everywhere there is a y (or whatever variable you're not differentiating w/respect to)

OpenStudy (binary3i):

by the way the answer is 1/10

OpenStudy (anonymous):

ah okay so let me try this again

OpenStudy (anonymous):

and 1/10 isn't an answer choice :P oh calculus...

OpenStudy (turingtest):

i got 10/3

OpenStudy (binary3i):

ok i will check it again

OpenStudy (turingtest):

is that a choice?

OpenStudy (anonymous):

the answer choices are 5/2, 4/3, 5/12, -7/6, and 2/3

OpenStudy (turingtest):

oh sorry I get 5/2 my bad... arithmetic :P

OpenStudy (anonymous):

lol np thanks! but i know for a fact ill see this one my exam and unfortunately i won't have OpenStudy with me >.<

OpenStudy (turingtest):

No, so you better try to see what you're missing so we can work on it. try the derivative of 3xy^2 again. This time keep the 3. Don't forget the product and chain rule.

OpenStudy (anonymous):

ah okay i got 6xy+3xy^2 but i can't figure out where to use the chain rule in this

OpenStudy (turingtest):

close again, but not quite... that x on the right should be gone, we took the derivative of it. plus we need the chain rule: when we take the derivative of y we need to multiply what we get by dy/dx wanna try again?

OpenStudy (anonymous):

differentiaion of 3xy^2 is read as .. 3 ( x into differentiation of y^2 + y^2 into differentiation of x).. now tell the answer :P

OpenStudy (turingtest):

here it is slow: d/dx(3xy^2)=3[d/dx(xy^2)]=3(y^2dx/dx+2xy*dy/dx) =3y^2+6xyy' where y' is the derivative dy/dx try to follow that step by step

OpenStudy (turingtest):

I can't see my LaTeX font right now so if somebody else wants to type it more clearly...

OpenStudy (anonymous):

what's the *?

OpenStudy (turingtest):

times

OpenStudy (turingtest):

d/dx(3xy^2)=3[d/dx(xy^2)]=3[y^2(dx/dx)+2xy(dy/dx)] =3y^2+6xyy'

OpenStudy (anonymous):

\[{d \over dx} 3xy^2 = 3 \left[ { d \over dx}(xy^2) \right] = 3 \left[ y^2 \left(dx \over dx \right) + 2xy \left( dy \over dx \right) \right] = 3y^2 + 6xyy''\]

OpenStudy (turingtest):

thanks^

OpenStudy (anonymous):

AH!!! NOW I GOT IT!

OpenStudy (turingtest):

cool, so what do you have for the derivative wrt x of 4x^(3)-3xy^(2)+y^(3)=28 ??? don't forget to take the derivative of everything, and use the chain rule ...and product rule.

OpenStudy (anonymous):

eeekk i don't think this is right at all. -12x^(2)/ [-3Y^(2)+6xy+3y^(2)]

OpenStudy (turingtest):

first of all, what happened to the other side of our equation?

OpenStudy (anonymous):

well i copied the way my calc teacher told us to do it (which he is the worst). dy/dx= -12x^(2)/ [-3Y^(2)+6xy+3y^(2)]

OpenStudy (turingtest):

??? that is madness, you should be doing it the way you see above as done by eashmore what is d/dx of... 1)4x^3 2)-3xy^2 3)y^3 take each individually Another thing, you forgot y' from the chain rule. That is critical to our problem. Whenever we are taking the derivative wrt x of y we need to wind up with a y'.

OpenStudy (anonymous):

i know... im one of the worst in the class... anyways, 1) 12x^2 2) 3y^2+6xyy' 3) 3y^2?

OpenStudy (turingtest):

good except for the last one... what is it missing?

OpenStudy (anonymous):

y'?

OpenStudy (turingtest):

right, because we took the derivative of a y term.

OpenStudy (turingtest):

oh sorry, actually you forgot that we made the second one negative as well 1)4x^3 1)12x^2 2)-3xy^2 2)-6xyy'-3y^2 3)y^3 3)3y^2y' so now that we have our derivatives we can add up those terms. what will be on the right side of our equation. Write the whole equation if you can.

OpenStudy (anonymous):

was i suppose to get this? 12x^2-3y^2-6xyy'+3y^2y'=0 ?

OpenStudy (turingtest):

yes, nice! now do you think you can solve it for y' ?

OpenStudy (anonymous):

ah! that's what i wrote on my paper!(except i had a plus 6 instead of minus) so after that is where i messed up, solving for y'

OpenStudy (turingtest):

perhaps, if you say so. try it again!

OpenStudy (anonymous):

okay so so far i have \[-6xyy'+3y ^{2}y'=-12x ^{2}+3y ^{2}\] is that a good start?

OpenStudy (turingtest):

good so far now factor out y' from the left...

OpenStudy (turingtest):

or note that you could divide everything by 3 first if you want to cut out a few numbers

OpenStudy (anonymous):

so then as the final thing \[y'= -4x ^{2}+y ^{2}\div -2xy+y ^{2}\]

OpenStudy (anonymous):

the division was suppose to be something over something but i didn't know how to do that

OpenStudy (turingtest):

I can't see that font right now anyway, but it looks like y'=(-4x^2+y^2)/(-2xy+y^2) right?

OpenStudy (anonymous):

ohh okay but yeah that's what it looks like

OpenStudy (turingtest):

So now we just plug in our point: (x,y)=(3,4)

OpenStudy (anonymous):

ah so that's how we will get 5/2?

OpenStudy (turingtest):

hopefully :)

OpenStudy (anonymous):

ah yay yay yay! :D thank you ever so much!!! <3

OpenStudy (turingtest):

no problem things to take away: 1)practice the product rule 2)practice the chain rule not that long of a list. Well congratulations, any questions?

OpenStudy (anonymous):

for this question no im find :) and yes im definitely going to have to practice those

OpenStudy (anonymous):

fine*

OpenStudy (turingtest):

well, see ya later, good luck :)

OpenStudy (anonymous):

see ya! thanks again! (: